### v3

@@ Abstract
Agentic large language model (LLM) training often involves multi-turn interaction trajectories that branch into multiple execution paths due to concurrent tool use, think-mode, sub-agent, context management and other runtime designs. As a result, the token produced by a single task naturally forms a tree-structured token trajectory with shared prefixes, rather than a linear sequence. Existing training pipelines linearize such trajectories and treat each branch independently, leading to substantial redundant computation in both forward and backward passes. To eliminate such redundancy, we introduce Tree Training, an efficient training framework for tree-structured trajectories. Its core component, Gradient Restoration, enables correct gradient aggregation across shared prefixes with negligible overhead, allowing each prefix to be computed exactly once while remaining mathematically equivalent to independent training on all branches. To support large trajectory trees in practice, we redesign the training engine to natively ingest tree-structured data and propose Tree Packing, a memory-efficient partitioning strategy that preserves high prefix reuse. Experiments conducted on dense and MOE models of real-world agentic trajectories show 6.2× training speedup for both supervised fine-tuning and the model update phase in reinforcement learning.
@@
Figure 1: Illustration of shared prefixes. Illustration of an example of a thinking model performing a multi-turn task: three sequences correspond to the model inputs and responses in the 1st, 2nd, and 3rd turns, respectively. In each new turn, the thinking process from the previous turn is discarded. This results in a situation where the concatenation of the previous turn’s input and output is not equivalent to the context fed into the current turn. As a result, upon completion of this multi-turn interactive task, it is impossible to represent all generated tokens for the task using a single tensor. Consequently, training with multiple sequences that share large overlapping prefixes leads to significant redundant computation.
@@ 1 Introduction
Training Large Language Models (LLMs) in agentic scenarios involves multi-turn interactions with environments, where a single task frequently branches into multiple parallel paths. This branching is driven by tree-structured workflow designs (e.g., tree structured planning Yao et al. (2023); Zhang et al. (2024) and sub-agentsTeam et al. (2026)), concurrent tool invocations Kim et al. (2024); Abdelaziz et al. (2024), context management Xu et al. (2025); Packer et al. (2024); Zhong et al. (2023), as well as runtime behaviors such as think-mode (discarding/replacing intermediate reasoning tokens between turns, e.g. Figure 1), and retokenization drift Luo et al. (2025). These mechanisms imply that the context fed into turn t+1t{+}1 is often not exactly the concatenation of the tokens produced in previous truns (from turn $1$ to turn $t$); therefore, after finishing a task, the collected tokens cannot always be organized as a single list/tensor, but instead form a tree-structured token trajectory with shared prefixes across branches. As a result, training must process multiple sequences with substantial shared prefixes, leading to significant redundant forward/backward computation.
@@
To address the above challenges, unlike caching-based approaches, we provide a strict mathematical derivation and develop an extremely low-overhead method, Gradient Restoration, to remove redundant prefix computation in the backward pass. Concretely, we apply a per-token compensation term so that each shared prefix is computed once while the resulting gradients are mathematically equivalent to training on all branches independently.
With this guarantee of efficient and correct prefix reuse, we further modify the training engine to ingest tree-structured trajectory data (where existing training frameworks Paszke et al. (2019); Abadi et al. (2016) typically only support list-like tensors). We also identify and address a practical issue for tree-structured training: a single trajectory tree can be too large to fit entirely within GPU memory, so we must partition a large tree into multiple subtrees and process them in multiple steps. The partitioning must ensure that each subtree fits in memory while preserving a high prefix-reuse ratio to achieve near-optimal compute efficiency. We implement Tree Packing with a heuristic DFS partitioning strategy, which is highly effective in practice and yields performance that closely approaches the theoretical limit.
Together, Gradient Restoration, Tree Packing, and a tree-aware training-engine redesign form Tree Training: an efficient framework for training LLMs on arbitrary tree-structured trajectories for both SFT and RL (reinforcement learning), achieving up to 6.2× end-to-end training speedup in our experiments.
@@
(2) We introduce Tree Training, a paradigm that executes each shared prefix once while correctly aggregating activations and gradients across descendant branches through Gradient Restoration, Tree Packing, and the redesign of the training engine.
@@ 2 Related Work
Sequence packing Krell et al. (2021) refers to concatenating multiple short sequences into a fixed-length training example to improve training efficiency. Our proposed tree packing method generalizes this idea: it not only concatenates sequences, but also introduces a strategy for efficiently segmenting overly long training data structured as a prefix tree (noting that a sequence is a special case of a prefix tree).
@@
Figure 2: Schematic of the data preprocess, forward pass, and backward pass in a tree-structured dataset. Pink blocks represent the prefix, yellow blocks correspond to two suffixes. The $(Q_{1},K_{1},V_{1},O_{1},dO_{1},dV_{1})$ correspond to the query, key, value, output, gradient of output and gradient of value matrices of the shared prefix between the two sequences. $S_{mn}$ denotes the result of the matmul between $Q_{m}$ and $K_{n}$, $P_{mn}$ denotes the result obtained by applying the softmax function to $S_{mn}$. In the forward pass, the outputs $O{{}_{1}}$ and $O{{}_{1}}{{}^{\prime}}$ are identical, thus making caching feasible. In the backward pass, the intermediate $dV{{}_{1}}$ and $dV{{}_{1}}{{}^{\prime}}$ are not identical, thus making caching unfeasible.
@@ 3 Tree Training
3.1 Overview
The standard approach for reusing computations of a shared prefix across two sequences is to employ prefix-caching. Although it is effective during inference and the forward pass of training, it can not be applied on the backward pass. Figure 2 showed this asymmetry in forward and backward passes of the attention operator.
Forward Passes
The forward pass of attention is expressed as:
$\begin{gathered}S=Q\times K^{T}\\ P=softmax(S)\\ O=P\times V\end{gathered}$ (1)
Due to the causal masking mechanism, the attention matrix $P$ is a Lower Triangular Matrix. The output $O_{i}$ for the $i$-th token depends solely on the token itself and its preceding tokens ($j\leq i$). Consequently, identical prefixes across different sequences share identical $P$ and $V$ matrices for the prefix positions, guaranteeing that their outputs are identical (i.e., $O_{1}\equiv O^{\prime}_{1}$ in Figure 2). This enables caching feasible and effective.
Backward Passes
The backward pass of $V$ is expressed as:
$\begin{gathered}dV=P^{T}\times dO\end{gathered}$ (2)
The transposition transforms the matrix $P$ into $P^{T}$, which is an Upper Triangular Matrix and is anti-causal. This implies that the gradient $dV_{i}$ for the $i$-th token aggregates gradients from the token itself and all subsequent tokens ($j\geq i$). Even if two sequences share an identical prefix, their differing suffixes introduce distinct $dO_{suffix}$, which propagate backward to the prefix. As a result, the gradients for the identical prefix positions ($dV_{prefix}$) will diverge (i.e., $dV{{}_{1}}$ and $dV{{}_{1}}{{}^{\prime}}$ in Figure 2).
While the forward pass is causal (prefix affects suffix), the backward pass is anti-causal (suffix affects prefix). Caching intermediate states for the backward pass is therefore not feasible across different suffixes, as the gradient computation for the prefix is inextricably coupled with the specific content of the entire suffix, making such caching memory-inefficient. To address these, we propose gradient restoration method in Section 3.2. Although our solution eliminates redundant computation, it needs a new data packing algorithm for large trees, which will be introduced in Section 3.3.
3.2 Gradient Restoration
Figure 3: A Comparative Illustration of Backward V-Gradient Computation and the Baseline. Pink blocks represent the $(Q,K,dO,dV)$ for the prefix parts, while yellow blocks correspond to those of the suffix parts. The orange blocks signify that their corresponding (S or P) values are active in the attention computation, whereas the white blocks indicate that their (S or P) values are masked out, the green blocks represent the (S or P) computations that can be omitted in our method.
Figure 4: Illustration of Tree Packing under memory constraints. Ideally, tree packing merges all trajectories into a single sequence to maximize prefix sharing. However, when the total tree size (e.g., 83k tokens) exceeds the GPU memory limit (e.g., 60k tokens), our method partitions the tree into minimal sub-sequences. Compared to the baseline flattening approach (164k tokens), our method (102k tokens) significantly reduces redundancy.
Figure 5: Real agentic trajectory trees and their overlap characteristics. The upper row visualizes representative trajectory trees extracted from multi-turn agentic RL rollouts, exhibiting different degrees of prefix sharing: Low Overlap, Medium Overlap, and High Overlap. The lower row plots the corresponding active trajectory counts over trajectory length for both the baseline (dashed) and Tree Training (solid). The area under each curve represents the total number of tokens processed by each method, and the ratio between the baseline and Tree Training areas reflects the theoretical token reuse achievable through tree-based computation sharing.
During the backward pass, the gradients of the prefix tokens are no longer identical across different sequences. The primary challenge, therefore, lies in determining how to eliminate gradient discrepancies such that the shared prefix can be computed only once across packed sequences, while preserving computational correctness.
We next contrast the baseline and our method in terms of backpropagation, and derive a gradient restoration algorithm that compensates for the gradient loss introduced by omitting redundant computations on shared prefixes.
3.2.1 baseline method
The baseline aligns with the current standard practice for training on branched trajectories, as most existing training frameworks only support taking batches of list-like tensors as training inputs. Specifically, for each leaf node in a trajectory tree, we construct a sequence by tracing back all tokens along the path from the leaf to the root; each leaf node thus corresponds to one sequence. These sequences are then processed through forward and backward passes individually, and the gradients accumulated across them are used to compute the overall gradient corresponding to the trajectory tree. However, this approach repeatedly recomputes shared prefixes across branches, leading to excessive redundant computation and thus an impractically long training time.
3.2.2 Parameter Updates with Gradients
We use a recursive definition over the trajectory tree. Here $i$ is the index of a tree node; $\operatorname{token}(i)$ denotes the token sequence represented by node $i$; $\operatorname{child}(j,i)$ denotes the index of the $j$-th child of node $i$ (and returns empty if node $i$ is a leaf node); and $X(i)$ denotes the list obtained by serializing the subtree rooted at node $i$.
The baseline serializes the subtree by pairing the prefix tokens $\operatorname{token}(i)$ with each child subtree, and concatenating these prefix–child pairs:
@@ 3.2.2 Parameter Updates with Gradients
$X_{base}(i)=\operatorname{Concat}[$ $\operatorname{token}(i),X_{base}(\operatorname{child}(0,i)),$ (3)
@@
$\operatorname{token}(i),X_{base}(\operatorname{child}(1,i)),\dots,$
$\operatorname{token}(i),X_{base}(\operatorname{child}(m,i))]$
Our method serializes the subtree by concatenating $\operatorname{token}(i)$ only once and reusing it across all children:
$X_{ours}(i)=\operatorname{Concat}[$ $\operatorname{token}(i),X_{ours}(\operatorname{child}(0,i)),$ (4)
$X_{ours}(\operatorname{child}(1,i)),\dots,X_{ours}(\operatorname{child}(m,i))]$
It suffices to show that performing backward on $X_{ours}(i)$ yields gradients equivalent to performing backward on $X_{base}(i)$ for the 2-level subtree consisting of node $i$ and its children; by recursion, this guarantees equivalence over the entire tree.
Therefore, we can focus on a simplified 2-level case. To avoid an abuse of notation, we let $X$ denote the serialized list of a root node with $n$ children. For a linear transformation: $Y=X\times weight$. The gradient $d\,\text{weight}=X^{T}\times dY$.
In the baseline serialization, we can view the process as first forming $n$ prefix–suffix pairs, each consisting of the shared prefix $P$ and one suffix $S_{i}$, and then concatenating these pairs together. Concretely, the matrix $X$ serialized by the baseline method is:
$\begin{gathered}x_{i}=[P;S_{i}],\forall i\in[1,n]\\ X_{base}=[x_{1};x_{2};...;x_{n}]\end{gathered}$ (5)
In contrast, the matrix X obtained by our serialization method is:
$\begin{gathered}X_{ours}=[P;S_{1};S_{2};...;S_{n}]\end{gathered}$ (6)
where P denotes prefix token, S denotes suffix token.
The computation of $dweight=X^{T}*dY$ can be interpreted as each column of $X^{T}$ multiplied by each row of $dY$, which yields the contribution of $token_{i}$ to the entire weight gradient, denoted as $dweight_{i}$. The total dweightdweight is equal to the sum of the $dweight_{i}$ for each individual token. Hence, if we can ensure that:
$\begin{gathered}P^{T}\times dY_{P}^{ours}=P^{T}\times(\sum_{i=1}^{n}dY^{base}_{p_{i}})\end{gathered}$ (7)
We can guarantee that the contribution of the prefix P to dweight in our method will be equivalent to its contribution in the baseline method. Equation 7 serves as the foundation for our gradient correction.
3.2.3 Algorithm and Analysis
In order to guarantee the equivalence of our Gradient Restoration method and the baseline method with respect to gradient updates, the two equations below must be identically satisfied for every trainable linear parameter.
$\begin{gathered}P^{T}\times dY_{P}^{ours}=P^{T}\times(\sum_{i=1}^{n}dY^{base}_{p_{i}})\\ \sum_{i=1}^{n}S_{i}^{T}\times dY_{S_{i}}^{ours}=\sum_{i=1}^{n}S_{i}^{T}\times dY_{S_{i}}^{base},\forall i\in[1,n]\end{gathered}$ (8)
Given that the values of P and S are determined exclusively by the dataset, which is identical for both serialization methods, the requirement reduces to ensuring that:
$\begin{gathered}dY_{P}^{ours}=dY^{base}_{p_{1}}+dY^{base}_{p_{2}}+...+dY^{base}_{p_{n}}\\ dY_{S_{i}}^{ours}=dY_{S_{i}}^{base},\forall i\in[1,n]\end{gathered}$ (9)
The remainder of this chapter will focus on explaining how to achieve identical satisfaction of the two aforementioned equations in the Transformer model through a straightforward correction.
Linear Operation
The linear operation is ubiquitous in the Transformer architecture. Assuming the following linear model:
$\begin{gathered}Y=X\times weight\end{gathered}$ (10)
The gradient for the input, $dX$, used during backpropagation is calculated as follows:
$\begin{gathered}dX=dY\times weight^{T}\end{gathered}$ (11)
The computation of $dX=dY\times weight^{T}$ can be viewed as each row of $dY$ being multiplied by the entire matrix $weight^{T}$ to yield the corresponding row of $dX$. This shows that a linear operation is pointwise: each $dX_{i}$ is obtained by multiplying $dY_{i}$ by $weight^{T}$, without depending on the gradients of preceding or subsequent tokens. Therefore, during backpropagation, if Equation 9 holds, then:
$\begin{gathered}dX_{P}^{ours}=dX^{base}_{p_{1}}+dX^{base}_{p_{2}}+...+dX^{base}_{p_{n}}\\ dX_{S_{i}}^{ours}=dX_{S_{i}}^{base},\forall i\in[1,n]\end{gathered}$ (12)
Equations 9 and 12 reveal that the gradient correction is transitive across linear operations. In the overall training pipeline, we only need to insert a gradient scaling step before the backward propagation. This operation restores the correct influence of each sequence’s prefix on the model update, ensuring training consistency. At the same time, for linear- type computations, we can avoid redundant computation on shared prefixes altogether.
Attention Operation
Assume $Q,K,V\in\mathbb{R}^{S\times d}$ are the inputs to the Attention Operation, where S denotes the sequence length and d is the head dimension. The forward pass and backward pass can be expressed as Equation 1 and Equation 2.
For the sake of clarity, this section will exemplify our gradient correction method for the attention operation using the computation of dV, which is illustrated in Figure 3.
Since the training dataset and weights (such as $W_{q}$ and $W_{k}$) are consistent between the two packing methods, the $S_{ij}$ and $P_{ij}$ at corresponding positions are equal.
According to the matrix computation depicted in Figure 3, if the following condition holds:
$\begin{gathered}dO_{P}^{ours}=dO^{base}_{p_{1}}+dO^{base}_{p_{2}}+...+dO^{base}_{p_{n}}\\ dO_{S_{i}}^{ours}=dO_{S_{i}}^{base},\forall i\in[1,n]\end{gathered}$ (13)
then:
$\begin{gathered}dV_{P}^{ours}=dV^{base}_{p_{1}}+dV^{base}_{p_{2}}+...+dV^{base}_{p_{n}}\\ dV_{S_{i}}^{ours}=dV_{S_{i}}^{base},\forall i\in[1,n]\end{gathered}$ (14)
This demonstrates that the gradient correction also exhibits transitivity within the attention operation, implying that the same correction method used for linear operations can be applied.
Other Operations
Taking the Rope operation as a case study, this section explains our general strategy for performing gradient corrections on some specific pointwise operations within the Transformer.
For any pointwise operation $Y_{i}=f(X_{i})$, the gradient with respect to the input, dX, in backpropagation can be calculated as follows:
$\begin{gathered}dX_{i}=dY_{i}*\frac{\partial Y_{i}}{\partial X_{i}}\end{gathered}$ (15)
To satisfy the transitivity of gradient correction outlined in Equations 12 and 9, the following condition should be guaranteed:
$\begin{gathered}\frac{\partial Y_{i}^{ours}}{\partial X_{i}^{ours}}=\frac{\partial Y_{i}^{base}}{\partial X_{i}^{base}}\end{gathered}$ (16)
For the Rope operation, defined as $Y=Rope(X,m)$ with position-dependent m, the gradient $\frac{\partial Y_{i}}{\partial X_{i}}$ depends on m. Consequently, it is imperative that the forward pass of Rope in our serialization method uses position_id values identical to those in the baseline method. Equation 16 decouples the entire backpropagation chain, allowing us to correct a specific operation without considering the influence of previous or subsequent operators.
3.2.4 Implementation
Building upon the analysis in 3.2.3, our implementation comprises three key components. Firstly, during the forward pass, we implement a shared prefix attention mask and correct position embedding by aligning each element with its original position in the original trajectory. Secondly, we introduce a Gradient Scaler mechanism in the backward pass to compensate for the gradient loss from skipped computations.
Figure 6: Implementation of flattened tree trajectory. Each flattened tree trajectory requires (1) a gradient scale tensor for prefix reuse, (2) a position embedding tensor that restores original token positions, and (3) a shared-prefix attention mask that enables proper computation reuse across overlapping prefixes during both forward and backward passes.
Shared Prefix Attention Mask
In the forward pass, due to tree packing, we introduce a shared prefix mask to ensure correctness in attention computation. The shared prefix mask is a modified causal mask that restricts the attention scope of each token so that tokens of different trajectories can safely share the same prefix representation without leaking information across branches. Based on the Flash Attention V3 project Shah et al. (2024); Wang et al. (2025), we have implemented a high-performance attention GPU kernel with node level shared prefix mask.
Position Embedding
The position_ids of the tree-packed data differs from those before packing. Therefore, we need to recompute the position IDs. In sequence packing, position_ids are simply reassigned based on the packed sequence layout. However, in tree training, we must instead restore the original (pre-packing) position IDs onto the packed sequences, so that each token’s positional information remains consistent with its original trajectory. A schematic diagram of the shared prefix mask and position embedding is shown in Figure 6.
Gradient Scaler
Based on the analysis in 3.2.3, we can safely recover the correct gradients by scaling. Specifically, after we calculate how many sequences reuse each node in the Tree — a coefficient we call the tree-scale, we multiply the gradient of each node by its corresponding tree-scale factor. The workflow of the Gradient Scaler is illustrated in Figure 7.
Figure 7: Backward computation with gradient scaling in Tree Training. After the first gradient of the flattened tree trajectory (dYdY) is computed, the gradient scaler is applied to scale the gradients of shared prefixes by their reuse counts. For example, the shared prefix $r\!\rightarrow\!u$ is used by five trajectories (scale = 5), and $v_{1}$ is used by three trajectories (scale = 3), ensuring correct gradient accumulation.
In the overall training pipeline, we only need to insert a gradient scaling step before the backward propagation. This operation restores the correct influence of each sequence’s prefix on the model update, ensuring training consistency. At the same time, for linear-type computations, we can avoid redundant computation on shared prefixes altogether.
Here, the gradient scaler acts as a correction factor that compensates for how many packed sequences share the same prefix. In other words, if a prefix node is reused by multiple sequences, its gradient is scaled accordingly so that the total gradient contribution remains mathematically equivalent to processing each sequence independently.
Parallelism in Distributed System
Tree Training is orthogonal to standard parallelism strategies (TP/EP/
DP/PP) and can be seamlessly combined with them. Context parallelism only affects self attention module. For the self-attention, different ranks receive distinct query shards along with the full key and value tensors per attention head. The attention masks for each parallel rank are generated based on their respective query shards, ensuring the correctness of the implementation.
MoE model load balancing
For models without auxiliary loss (e.g., Qwen3 MoE): The tree structure does not affect the load balancing process.
For models with auxiliary loss: We achieve mathematical equivalence to the baseline by multiplying each prefix token by its share count (the gradient scaler) during the router auxiliary loss calculation.
3.3 Tree Packing

### v4

@@ Abstract
Agentic large language model (LLM) training often involves multi-turn interaction trajectories that branch into multiple execution paths due to concurrent tool use, think-mode, sub-agent, context management and other runtime designs. As a result, the tokens produced by a single task naturally form a tree-structured token trajectory with shared prefixes, rather than a linear sequence. Existing training pipelines linearize such trajectories and treat each branch independently, leading to substantial redundant computation in both forward and backward passes. We derive that averaging the loss over all branches independently is algebraically identical to a per-token weighted loss, where each token’s weight equals the fraction of branches passing through it. The problem therefore reduces to computing the log-probability of every token in the prefix tree exactly once, with no repeated computation across shared prefixes: we propose DFS serialization of the tree, which visits every token exactly once, and adapt full-attention and SSM layers to ensure the resulting log-probabilities match independent per-branch calculation exactly. In practice, a single trajectory tree can be too large to fit in GPU memory; we therefore propose Tree Partitioning, a memory-efficient partitioning strategy that splits the tree into subtrees each fitting within GPU memory while preserving high prefix reuse. Together, these contributions form Tree Training, an efficient framework for training LLMs on tree-structured trajectories, achieving up to 6.2× end-to-end training speedup on dense and MOE models for both supervised fine-tuning and reinforcement learning.
@@
Figure 1: Trajectory tree and DFS sequence. Left: a trajectory tree with $K=3$ root-to-leaf paths. Each node $n$ holds a token segment $T_{n}$; $g_{n}$ is the number of paths through $n$ (blue nodes are shared prefixes with $g_{n}>1$). Node text in red is model output (trained); black is user/environment input. Right: the DFS sequence places all nodes in depth-first order so that every token is computed exactly once, with no redundancy.
@@ 1 Introduction
Training Large Language Models (LLMs) in agentic scenarios involves multi-turn interactions with environments, where a single task frequently branches into multiple parallel paths. This branching is driven by tree-structured workflow designs (e.g., tree structured planning Yao et al. (2023); Zhang et al. (2024) and sub-agentsTeam et al. (2026)), concurrent tool invocations Kim et al. (2024); Abdelaziz et al. (2024), context management Xu et al. (2025); Packer et al. (2024); Zhong et al. (2023), as well as runtime behaviors such as think-mode (discarding/replacing intermediate reasoning tokens between turns, e.g. Figure 1), and retokenization drift Luo et al. (2025). These mechanisms imply that the context fed into turn t+1t{+}1 is often not exactly the concatenation of the tokens produced in previous turns (from turn $1$ to turn $t$); therefore, after finishing a task, the collected tokens cannot always be organized as a single list/tensor, but instead form a tree-structured token trajectory with shared prefixes across branches. As a result, training must process multiple sequences with substantial shared prefixes, leading to significant redundant forward/backward computation.
@@
To address the above challenges, we mathematically derive that the baseline loss—averaging over all branches independently—is algebraically identical to a per-token weighted loss where each token’s weight equals the fraction of branches passing through it. The problem therefore reduces to computing the log-probability of every token in the prefix tree exactly once, with no repeated computation across shared prefixes. We propose DFS serialization of the tree, which visits every token exactly once, and adapt full-attention and SSM layers to ensure the resulting log-probabilities match the baseline exactly.
We further redesign the training engine to natively ingest tree-structured data. We also address a key practical constraint: a single trajectory tree is often too large to fit entirely within GPU memory. We therefore propose Tree Partitioning, a memory-efficient partitioning strategy that splits a large tree into subtrees each fitting within memory while preserving a high prefix-reuse ratio.
Together, DFS serialization with model-layer adaptations, Tree Partitioning, and the training-engine redesign form Tree Training: an efficient framework for training LLMs on tree-structured trajectories for both SFT and RL, achieving up to 6.2× end-to-end training speedup in our experiments.
@@
(2) We introduce Tree Training, a paradigm that executes each shared prefix once while achieving mathematically equivalent training to the baseline, via DFS serialization, targeted model-layer adaptations, Tree Partitioning, and the redesign of the training engine.
@@ 2 Related Work
Sequence packing Krell et al. (2021) refers to concatenating multiple short sequences into a fixed-length training example to improve training efficiency. Our proposed tree partitioning method generalizes this idea: it not only concatenates sequences, but also introduces a strategy for efficiently segmenting overly long training data structured as a prefix tree (noting that a sequence is a special case of a prefix tree).
Recent hybrid LLM architectures interleave full-attention layers with State Space Model (SSM) layers. Gated Delta Net (GDN) Yang et al. (2025c) is a recent SSM variant that improves over earlier SSM architectures such as Mamba Gu and Dao (2024) and Mamba2 Dao and Gu (2024) by incorporating the delta rule, achieving better associative recall and in-context learning. GDN is adopted as the SSM component in hybrid models such as Qwen3.5 Team (2026). Unlike full-attention layers, SSM layers maintain a recurrent hidden state that is passed sequentially between tokens, which requires dedicated handling when processing tree-structured data.
@@
Figure 2: SSM state routing. Sequential flow (left) feeds $n_{4}$’s state into $n_{2}$—wrong, since $n_{2}$ is a sibling of $n_{1}$, not a descendant of $n_{4}$. Tree routing (right) routes each chunk to its parent’s state: both $n_{3}$ and $n_{4}$ read $s_{n_{1}}$, while $n_{2}$ correctly reads $s_{n_{0}}$.
@@ 3 Tree Training
A trajectory tree is a rooted tree in which each node $n$ holds a token segment $T_{n}$. Each root-to-leaf path spells out a complete trajectory by concatenating the token segments of its nodes in order. Let $\mathcal{T}$ denote the set of all nodes and $\mathcal{P}=\{p_{1},\dots,p_{K}\}$ the set of $K$ root-to-leaf paths (one per leaf). An example with $K=3$ is shown in Figure 1. Goal. Process all tokens in the tree with each token computed exactly once (no redundant computation across shared prefixes) in a single forward and backward pass, and produce gradients that are mathematically equivalent to the baseline of running all $K$ paths independently and averaging their losses.
Derivation. The baseline sep-avg loss runs each of the $K$ paths independently:
$L_{\text{sep\_avg}}=\frac{1}{K}\sum_{k=1}^{K}\sum_{t\in p_{k}}\ell_{t}(\theta)$ (1)
where $\ell_{t}(\theta)$ is any per-token objective. For SFT, $\ell_{t}(\theta)=-\log p_{\theta}(y_{t}\mid x_{\leq t})$; for RL with policy gradient, $\ell_{t}(\theta)=-A_{t}\cdot\log p_{\theta}(y_{t}\mid x_{\leq t})$ where $A_{t}$ is the per-token advantage.
Eq. (1) iterates over paths then over tokens in each path. We swap the order: iterate over every unique token $t$ and count how many paths include it. Token $t$ belongs to node $n$; let $g_{n}=|\{p\in\mathcal{P}:n\in p\}|$ be the number of paths passing through $n$, and $g_{t}\triangleq g_{n}$ for $t\in T_{n}$. Token $t$ is therefore counted $g_{n}$ times in the double sum:
$\sum_{k=1}^{K}\sum_{t\in p_{k}}\ell_{t}(\theta)\;=\;\sum_{n\in\mathcal{T}}\sum_{t\in T_{n}}g_{n}\cdot\ell_{t}(\theta)\;=\;\sum_{t}g_{t}\cdot\ell_{t}(\theta)$ (2)
Substituting (2) into (1):
$L_{\text{sep\_avg}}=\sum_{t}\frac{g_{t}}{K}\cdot\ell_{t}(\theta)$ (3)
We therefore define the tree loss:
$\boxed{L_{\text{tree}}=\sum_{t}\frac{g_{t}}{K}\cdot\ell_{t}(\theta)}$ (4)
By linearity of differentiation:
$\frac{\partial L_{\text{tree}}}{\partial\theta}=\sum_{t}\frac{g_{t}}{K}\cdot\frac{\partial\ell_{t}}{\partial\theta}{=}\frac{\partial L_{\text{sep\_avg}}}{\partial\theta}\qquad$ (5)
This derivation is model-agnostic and objective-agnostic: since it relies only on the additive structure of $\ell_{t}$ across tokens and paths—not on its specific form—it applies to any per-token objective. The coefficient $g_{t}/K$ arises from equal weighting of $K$ paths ($w_{k}=1/K$); in the general case with path weights $\{w_{k}\}$, the per-token weight becomes $\sum_{k:\,t\in p_{k}}w_{k}$. For RL with policy gradient, $\ell_{t}(\theta)=-A_{t}\cdot\log p_{\theta}(y_{t}\mid x_{\leq t})$ with per-token advantage $A_{t}$; the $g_{t}/K$ factor remains the correct path-count normalization, while $A_{t}$ is already absorbed into $\ell_{t}$. Note that $L_{\text{tree}}$ and $L_{\text{sep\_avg}}$ are identical expressions—the substitution is purely algebraic; it does not yet impose any constraint on how $\ell_{t}(\theta)$ is computed in practice.
Problem reduction. The goal therefore reduces to: find a method that (1) computes each token exactly once with no redundant computation, and (2) ensures that $\ell_{t}(\theta)$ computed this way matches what it would be in a standalone per-branch forward pass—i.e., forward equivalence:
$\ell_{t}^{\text{tree}}(\theta)=\ell_{t}^{p}(\theta)\quad\forall\,p\in\mathrm{paths}(n_{t})$ (6)
We achieve this via DFS traversal of the token prefix tree, which serializes all nodes in depth-first order so each token appears exactly once, replacing the baseline serialization that repeats prefixes:
@@
$X_{\text{base}}(i)=\operatorname{Concat}\bigl[\operatorname{token}(i),\,X_{\text{base}}(\operatorname{child}(0,i)),\\ \operatorname{token}(i),\,X_{\text{base}}(\operatorname{child}(1,i)),\;\dots\bigr]$ (7)
with:
@@
$X_{\text{DFS}}(i)=\operatorname{Concat}\bigl[\operatorname{token}(i),\,X_{\text{DFS}}(\operatorname{child}(0,i)),\\ X_{\text{DFS}}(\operatorname{child}(1,i)),\;\dots\bigr]$ (8)
How to enforce (6) on the DFS-serialized sequence—i.e., ensuring each token’s log-probability matches its per-branch value despite the reordering introduced by DFS traversal—is described in Section 3.1.
3.1 Implementation
Loss Weights.
For each token $t$ in the DFS-serialized sequence, its per-token loss is weighted by $g_{t}/K$ before reduction, where $g_{t}$ is the number of root-to-leaf paths passing through token $t$’s node and $K$ is the total number of leaves. This is applied as a single element-wise multiplication on the per-token loss tensor—requiring no modification to the backward pass and negligible overhead.
With the following two adaptations, the forward equivalence condition (6) is satisfied and the gradient equivalence holds.
Attention Mask.
We construct a tree attention mask such that token $i$ attends to token $j$ iff (1) $j\leq i$ and (2) the nodes of $i$ and $j$ lie on a common root-to-leaf path. This prevents cross-branch attention. We implement a GPU kernel based on Flash Attention V3 Shah et al. (2024); Wang et al. (2025) with node-level shared-prefix masking. The attention pattern is illustrated in Figure 3.
Figure 3: Tree attention mask. Left: the $6\times 6$ attention matrix for the DFS sequence; a cell is filled iff the query token can attend to the key token. Tokens from different branches (e.g. $n_{3}$ and $n_{2}$) are mutually blocked. Right: the corresponding tree structure; the dashed red arrow shows a blocked cross-branch path.
Position Embedding.
Each token must receive the same position ID as it would in its standalone per-path sequence. A token at offset $j$ within node $n$ is assigned:
$\mathrm{pos}(t)=\sum_{n^{\prime}\in\mathrm{anc}(n)\setminus\{n\}}|T_{n^{\prime}}|+j$ (9)
where $\mathrm{anc}(n)$ denotes the ancestors of $n$. Sibling nodes at the same depth share the same position range, matching the per-path baseline and ensuring RoPE Su et al. (2023) produces identical results.
SSM Layers (GDN).
For hybrid models that interleave full-attention with Gated Delta Net (GDN) Yang et al. (2025c) SSM layers, the attention mask alone is insufficient: GDN layers maintain a recurrent hidden state passed sequentially between chunk boundaries, so a naïve DFS layout would carry the state from one branch into the next sibling branch. We apply two fixes. We treat each tree node’s token segment as one chunk—the unit of SSM state transfer. (i) State routing: each chunk $c$ reads its initial SSM state from its parent chunk $\pi(c)$ rather than the preceding chunk in DFS order:
$\mathbf{h}_{c}^{(0)}=\mathbf{h}_{\pi(c)}^{(\mathrm{end})},\qquad\mathbf{h}_{-1}^{(\mathrm{end})}\triangleq\mathbf{0}$ (10)
DFS order guarantees $\pi(c)<c$, so parent states are always ready. Sibling chunks share the same parent state tensor, so autograd accumulates their gradients there automatically. Figure 2 illustrates the difference between sequential and tree state routing.
(ii) Causal convolution: SSM layers include a causal conv1d with kernel size $K_{\mathrm{conv}}$, which requires the preceding $K_{\mathrm{conv}}{-}1$ tokens as left context. In the DFS sequence, the tokens immediately before a child chunk belong to a sibling branch rather than the ancestor path, so using them as context is incorrect. Instead, each chunk $c$ uses $\mathbf{ctx}_{\pi(c)}$—the last $K_{\mathrm{conv}}{-}1$ output tokens of its parent chunk $\pi(c)$, saved after processing the parent—as its left context:
$\mathbf{y}_{c}=\mathrm{Conv}\!\left([\,\mathbf{ctx}_{\pi(c)}\,;\,\mathbf{x}_{c}\,]\right)_{[K_{\mathrm{conv}}-1:]}$ (11)
This replicates the context each token would see in an independent per-branch forward pass. Figure 4 illustrates this for $K_{\mathrm{conv}}=3$.
Figure 4: SSM causal convolution ($K_{\mathrm{conv}}=3$). When processing $n_{2}$ (child of $n_{0}$), the correct conv window is formed by prepending conv_state[$n_{0}$]—the stored last $K_{\mathrm{conv}}{-}1$ effective tokens of the parent—as left context, bypassing the DFS-adjacent sibling tokens ($n_{4}$) that occupy the preceding positions.
3.2 Tree Partitioning
Figure 5: Tree Partitioning under memory constraints. When the full tree (83k tokens) exceeds GPU memory (60k), our method partitions it into two sub-sequences (102k total) vs. the baseline’s 164k.